# LUME: LLM Unlearning with Multitask Evaluations

### v1

@@ 2.1 Benchmark Construction
Task 2 (Synthetic biographies with sensitive PII): We use rule based heuristics to generate personal biographies with following PII fields: a randomly generated name, a birthday randomly sampled between 01/01/1964 and 01/01/1991, a fake Social Security number (SSN) within the range 900-xx-xxxx (which can never belong to a real person ssa (2011)), a random phone number, an email address of the form firstname_lastname@me.com and a non-existent physical home addresses obtained by combining a random street address from a US state with an alternate city and zip-code from a different state. For each synthetic individual, we prompt the Mixtral model to create a short biography by including the fictitious PII information.

### v2

@@ 2.1 Benchmark Construction
Task 2 (Synthetic biographies with sensitive PII):
We use rule based heuristics to generate personal biographies with following PII fields: a randomly generated name, a birthday randomly sampled between 01/01/1964 and 01/01/1991, a fake Social Security number (SSN) within the range 900-xx-xxxx (which can never belong to a real person ssa (2011)), a random phone number, an email address of the form firstname_lastname@me.com and a non-existent physical home addresses obtained by combining a random street address from a US state with an alternate city and zip-code from a different state. For each synthetic individual, we prompt the Mixtral model to create a short biography by including the fictitious PII information.